40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Entry  Fri Jan 20 17:20:20 2017, Gabriele, Optics, Daily Progress, Input part of the optical lever setup installed and aligned optical_setup.png
    Reply  Mon Jan 23 16:29:52 2017, Gabriele, Optics, Daily Progress, Input part of the optical lever setup installed and aligned 
Message ID: 272     Entry time: Mon Jan 23 16:29:52 2017     In reply to: 270
Author: Gabriele 
Type: Optics 
Category: Daily Progress 
Subject: Input part of the optical lever setup installed and aligned 

Today I cabled and installed the four QPD, in a temporary position. I also assembled four picomotor mounts that will be used for the auto-centering.

Quote:

Today I installed and aligned part of the optical components for the optical lever of the new setup. For the moment being I installed only the input components, and aligned the beams into the vacuum chamber. Since I don't have any in-vacuum optics yet, there's nothing more that can be done now.

 

ELOG V3.1.3-